OpenStudy (anonymous):

Which is a solution to the equation 5x + 2y = –1?

OpenStudy (anonymous):

First find either x or y

OpenStudy (anonymous):

5x+2y=-1 -5x

OpenStudy (anonymous):

2y=-5x-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Divide by 2: y=-5/2x-1/2

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

u have answer choices? it says choose

OpenStudy (anonymous):

Is this system of equations? Should be 2 equations

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

5x - 2y = -1 x + 4y = 35?

OpenStudy (anonymous):

That makes sense then. those 2?

OpenStudy (anonymous):

A. (2, –4) B. (–1, 4) C. (–2, 5) D. (1, –3)

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

ok then... i guess u just plug in the points

OpenStudy (anonymous):

wat u mean?, :/

OpenStudy (anonymous):

A. (2, –4) B. (–1, 4) C. (–2, 5) D. (1, –3)

OpenStudy (anonymous):

let me show u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5x+2y=-1 Try A 5(2)+2(-4)=-1 10-8=-1 This is not the answer

OpenStudy (anonymous):

Try B

OpenStudy (anonymous):

5x+2y=-1 Try A 5(-1)+2(4)=-1 -5+8=-3

OpenStudy (anonymous):

Not the answer either

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

5x+2y=-1 5(-2)+2(5)=-1 -10+10=0 Nope D (1,-3) 5x+2y=-1 5(1)+2(-3)=-1 5-6=-1 The correct answer is D

OpenStudy (anonymous):

Do u get all of this?

OpenStudy (anonymous):

can we do some more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

just In case

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

What is the slope of the graph of 14x + 7y = –9?

OpenStudy (anonymous):

A. B. C. D.

OpenStudy (anonymous):

U just need to put it in slope intercept form (y=mx+b)

OpenStudy (anonymous):

A. -2 B.- 1/2 C. 1/2 D. 2

OpenStudy (anonymous):

14x + 7y = –9 -14x 7y=-14x-9 Divide by 7 y=-2x-9/7 m is the slope so

OpenStudy (anonymous):

U know the answer...

OpenStudy (anonymous):

no @Haiqa

OpenStudy (anonymous):

A. -2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Sorry bout that

OpenStudy (anonymous):

y=-2x-9/7 is the equation. remember slope-intercept form id y=mx+b

OpenStudy (anonymous):

-2 is in the place of m, i.e the slope

OpenStudy (anonymous):

That help?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wat about this one

OpenStudy (anonymous):

Plz ask a new question.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Its all crowded here. Thnx

OpenStudy (anonymous):

ok